A recipe asks for 8/9 cup of granola. How much is needed to make 7/5 of a recipe?
Answer by TimothyLamb(4379)

g = 8/9 * 7/5
g = (8*7)/(9*5)
g = 56/45
g ~= 1.2444 cups
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php